Three students have each saved $1,000. Each has an investment opportunity in which he or she can invest up to $2,000. Here are the rates of return on the students’ investment projects:

Student Return(Percent)

Shen 4

Antonio 7

Caroline 15

Assume borrowing and lending is prohibited, so each student uses only personal saving to finance his or her own investment project.

Complete the following table with how much each student will have a year later when the project pays its return.

Student Money a Year Later(Dollars)

Shen (....)

Antonio (....)

Caroline (....)


Now suppose their school opens up a market for loanable funds in which students can borrow and lend among themselves at an interest rate r

A student would choose to be a lender in this market if his or her expected rate of return is (...) than r


Suppose the interest rate is 6 percent.


Among these three students, the quantity of loanable funds supplied would be, (....) and quantity demanded would be. (....)


Now suppose the interest rate is 12 percent.


Among these three students, the quantity of loanable funds supplied would be, (....) and quantity demanded would be (....).

At an interest rate of (..%), the loanable funds market among these three students would be in equilibrium. At this interest rate (....),    would want to borrow, and  (....)  would want to lend.

Suppose the interest rate is at the equilibrium rate.


Complete the following table with how much each student will have a year later after the investment projects pay their return and loans have been repaid.

Student Money a Year Later(Dollars)

Shen (....)

Antonio (....)

Caroline (....)


True or False: Only borrowers are made better off, and lenders are made worse off.



1
Expert's answer
2021-02-18T13:35:24-0500

Student Money a Year Later (Dollars)

Shen

1040

Antonio

1070

Caroline

1150

 

Shen 1.04*1000= $1040

Antonio 1.07*1000= $1070

Caroline 1.15*1000 =$1150

A student would choose to be a lender in this market if his or her expected rate of return is (greater) than r.

Suppose the interest rate is 6 percent. Among these three students, the quantity of loanable funds supplied would be, ($1000) and quantity demanded would be. ($2000)

Now suppose the interest rate is 12 percent. Among these three students, the quantity of loanable funds supplied would be, ($2000) and quantity demanded would be ($1000).

At an interest rate of (7%), the loanable funds market among these three students would be in equilibrium. At this interest rate (Caroline), would want to borrow, and (Shen) would want to lend.

Student Money a Year Later (Dollars)

Shen

1070

Antonio

1070

Caroline

1230

 

Shen 1.07*1000= $1070

Antonio 1.07*1000= $1070

Caroline 1.15*2000 – 1.07*1000 =$1230

False.